How many quarts of water must be added to 40 quarts of solution which is 5 percent acid solution to dilute it to a 2 percent solution?
Thanks!

Answer by lwsshak3(11628)

How many quarts of water must be added to 40 quarts of solution which is 5 percent acid solution to dilute it to a 2 percent solution?
***
let x=amt of water to be added
40+x=amt of final mixture
..
5%*40+0%x=2%(40+x)
2+0=0.8+.02x
.02x=1.2
x=60
amt of water to be added=60 quarts